the perimeter of a retangle is 150 meters. The length is four times the width. What is the area of the retangle?
Answer by checkley77(12844)

L=4W
2L+2W=150
2*4W+2W=150
8W+2W=150
10W=150
W=150/10
W=15 ANS. FOR THE WIDTH.
L=4*15=60 ANS. FOR THE LENGTH.
PROOF:
2*60+2*15=150
120+30=150
150=150